# Zeroth-Order Constrained Optimization from a Control Perspective via Feedback Linearization

### v1

@@ Abstract
Designing safe derivative-free optimization algorithms under unknown constraints is a fundamental challenge in modern learning and control. Most existing zeroth-order (ZO) approaches typically assume white-box constraints or focus on convex settings, leaving the general case of nonconvex optimization with black-box constraints largely open. We propose a control-theoretic framework for ZO constrained optimization that enforces feasibility without relying on solving costly convex subproblems. Leveraging feedback linearization, we introduce a family of ZO feedback linearization (ZOFL) algorithms applicable to both equality and inequality constraints. Our method requires only noisy, sample-based gradient estimates yet provably guarantees constraint satisfaction under mild regularity conditions. We establish finite-time bounds on constraint violation and further present a midpoint discretization variant that further improves feasibility without sacrificing optimality. Empirical results demonstrate that ZOFL consistently outperforms standard ZO baselines, achieving competitive objective values while maintaining feasibility.
@@ 1 Introduction
Among derivative-free approaches, zeroth-order methods have attracted significant attention due to their simplicity and scalability to high dimensions Wang et al. (2018); Nesterov and Spokoiny (2017). The core idea is to build stochastic gradient estimators via finite differences of function evaluations and then plug them into standard gradient-based updates Berahas et al. (2022). For instance, two-point estimators perturb the decision along random isotropic directions (Gaussian or uniform on the unit sphere) and combine the function values to approximate gradients Nesterov and Spokoiny (2017); Tang et al. (2023); Ren et al. (2023). When combined with gradient descent, such estimators yield provable converge rates for unconstrained optimization.
@@
In the constrained setting, most existing ZO algorithms assume black-box objectives but white-box constraints. Explicit knowledge of the constraint set enables efficient projections or local linearizations, ensuring feasibility. This has led to a variety of algorithms, including projection–gradient-descent Tang et al. (2023); Liu et al. (2018a); Chen et al. (2019); Ghadimi et al. (2016), Frank–Wolfe–type methods Sahu and Kar (2020); Liu et al. (2024); Balasubramanian and Ghadimi (2019), and Sequential Quadratic Programming (SQP)-style approaches Curtis et al. (2023); Berahas et al. (2025). However, in many safe learning settings—such as safe RL or chance-constrained optimization (Achiam et al., 2017; Cui et al., 2020; Zhang and Li, 2011)—the constraint functions themselves are unknown. Here, only noisy zeroth-order estimates of the constraint gradients are available, and feasibility becomes much harder to enforce. Most existing work in this regime focuses on convex problems and relies on primal–dual schemes (Zhou et al., 2025; Chen et al., 2022a; Yi et al., 2021; Nguyen and Balasubramanian, 2022; Li et al., 2022b; Maheshwari et al., 2022; Liu et al., 2020; Chen et al., 2023; Hu et al., 2024). For nonconvex problems, guarantees are scarce: some works provide only empirical evidence (Zhou et al., 2025), while others require solving expensive convex subproblems at each step (Nguyen and Balasubramanian, 2022; Li et al., 2022b). Feasibility in these settings remains fragile, typically degrading as noise in the gradient estimators increases (Oztoprak et al., 2021).
@@
In contrast, in the first-order (FO) setting, SQP methods are known to be highly effective for nonconvex constrained optimization. They often outperform primal–dual approaches in practice, particularly when the number of constraints is small relative to the dimension of the decision variable (Gould et al., 2003; Liu et al., 2018b; Zhang and Luo, 2020; Applegate et al., 2024). Another complementary line of work interprets constrained optimization through a control-theoretic lens (Cerone et al., 2024; Zhang et al., 2025). Here, the optimization dynamics are viewed as a controlled system: the primal variables are states, the Lagrange multipliers are control inputs, and feasibility corresponds to stabilizing the system at an equilibrium. This perspective enables the use of nonlinear control tools, such as feedback linearization (FL), to design algorithms. Recent work shows that, under suitable conditions, FL-based schemes recover SQP-like updates and achieve strong performance on nonconvex problems (Zhang et al., 2025).
@@
Gap and motivation. Despite their promise in first-order optimization, FL and SQP approaches have not been systematically studied in the zeroth-order regime. Extending them is nontrivial: FL and SQP depends critically on precise first-order information, but in the zeroth-order setting only noisy estimates are available, breaking the mechanisms that ensure feasibility. This raises the fundamental question:
@@
Our Contributions. We develop a zeroth-order constrained optimization framework that extends feedback linearization ideas to the derivative-free regime, inspired by control and dynamical systems perspective (Cerone et al., 2024; Zhang et al., 2025). First, we show how to construct an FL scheme tailored to dynamics evolving under noisy gradient information, in contrast to prior approaches that rely on convex relaxations or primal–dual surrogates. Second, we demonstrate that full Jacobians are unnecessary: it is enough to approximate a small set of Jacobian–vector products, which can be efficiently estimated via two-point zeroth-order queries. Third, we establish theoretical guarantees (Theorem 1 and Theorem 2) showing that constraint violations contract toward zero with high probability, up to controllable approximation and discretization errors. Finally, we provide empirical evidence that our method consistently achieves stronger feasibility performance than standard baselines, while achieving competitive objective values.
@@ 2 Preliminaries
Problem setup.  We consider constrained optimization problems of the form
$\min_{x\in\mathbb{R}^{n}}f(x)\quad\text{s.t. }h(x)=0,$ (1)
@@
$\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)=0.$ (2)
@@ 2.1 First-order Constrained Optimization: A Control Perspective
Recent works Cerone et al. (2024); Zhang et al. (2025) interpret constrained optimization from a control perspective, offering new insights and enabling novel algorithmic designs in the first-order optimization regime. The key idea is to reinterpret Eq. 2 as the equilibrium of a dynamical system. Specifically, define the updates
$x_{t+1}-x_{t}=-\eta_{t}\!\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right)\!,~~~y_{t}=h(x_{t}),$ (3)
where $x_{t}$ is the system state, $y_{t}$ the constraint output, and $\lambda_{t}$ the control input.
@@
$x_{t+1}-x_{t}=F(x_{t})+G(x_{t})\lambda_{t},$ (4)
by introducing a new input that cancels the nonlinearities Isidori (1985); Henson and Seborg (1997). If $G(x)$ is invertible, one may set $\lambda_{t}=G(x_{t})^{-1}(u_{t}-F(x_{t}))$, so that the dynamics reduce to $x_{t+1}-x_{t}=u_{t}$, a linear system for which standard stabilizing controllers are available.
@@
$y_{t+1}-y_{t}\approx J_{h}(x_{t})(x_{t+1}-x_{t})=-\eta_{t}J_{h}(x_{t})\nabla f(x_{t})-\eta_{t}J_{h}(x_{t})J_{h}(x_{t})^{\top}\lambda_{t},$ (5)
where the terms can be viewed as $F(x_{t})$ and $G(x_{t})\lambda_{t}$ in Eq. 4. Choosing
$\lambda_{t}=-\big(J_{h}(x_{t})J_{h}(x_{t})^{\top}\big)^{-1}\!\left(J_{h}(x_{t})\nabla f(x_{t})-Kh(x_{t})\right),$
cancels the nonlinear dependence and yields the linearized dynamics $y_{t+1}-y_{t}\approx-\eta_{t}Kh(x_{t})=-\eta_{t}Ky_{t}$. By picking $K$ Hurwitz, the constraints converge exponentially to zero.
This design gives the first-order feedback linearization (FO-FL) method:
FO-FL (Equality Constraints) Cerone et al. (2024); Zhang et al. (2025) $\begin{split}x_{t+1}-x_{t}&=-\eta_{t}\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\\ \lambda_{t}&=-\left(J_{h}(x_{t})J_{h}(x_{t})^{\top}\right)^{-1}\!\left(J_{h}(x_{t})\nabla f(x_{t})-Kh(x_{t})\right).\end{split}$ (6)
FO-FL has been shown to effectively handle nonlinear dynamics, making it well-suited for nonconvex constrained optimization. Empirical and theoretical studies Cerone et al. (2024); Schropp and Singer (2000); Zhang et al. (2025) confirm its strong performance relative to primal–dual methods.
@@ 2.2 Zeroth-order Constrained Optimization: Baseline and Challenges
The absence of first-order information motivates the use of stochastic finite-difference estimators for $\nabla f(x)$ and $J_{h}(x)$. A standard choice is the two-point estimator:
$\begin{split}\widetilde{\nabla}f(x_{t})&=\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}\frac{f(x_{t}+r_{1}u_{i})-f(x_{t}-r_{1}u_{i})}{2r_{1}}\,u_{i},\\ \widetilde{J}_{h}(x_{t})&=\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}\frac{h(x_{t}+r_{1}u_{i})-h(x_{t}-r_{1}u_{i})}{2r_{1}}\,u_{i}^{\top},\end{split}$ (7)
@@
ZO-baseline for Equality-Constrained Optimization Cerone et al. (2024) $\begin{split}x_{t+1}-x_{t}&=-\eta_{t}\!\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right),\\ \lambda_{t}&=-\!\left(\widetilde{J}_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\!\!\left(\widetilde{J}_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\right).\end{split}$ (8)
This approach has been explored in recent work on noisy or biased estimators Oztoprak et al. (2021). However, it suffers from a critical drawback: constraint satisfaction is no longer guaranteed. To see this, note that the constraint dynamics become
$\textstyle h(x_{t+1})-h(x_{t})$ $\textstyle\approx J_{h}(x_{t})(x_{t+1}-x_{t})$
@@ 3.1 Algorithm
$\textstyle h(x_{t+1})-h(x_{t})\approx J_{h}(x_{t})(x_{t+1}-x_{t})=-\eta_{t}\!\left(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right).$ (9)
@@
$\textstyle\lambda_{t}=-\!\left(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\!\big(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\big),$ (10)
@@
Insight. A closer examination reveals that the full knowledge of $J_{h}(x_{t})$ is in fact unnecessary: we only need the products $J_{h}(x_{t})\widetilde{\nabla}f(x_{t})$ and $J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}$. In other words, instead of recovering the entire Jacobian, it suffices to evaluate its inner product along $m+1$ prescribed directions: the estimated gradient $\widetilde{\nabla}f(x_{t})$ and the rows of $\widetilde{J}_{h}(x_{t})$. Crucially, these Jacobian–vector products can be efficiently approximated using standard two-point estimators as follows, thereby rendering the scheme implementable in the ZO setting:
$\begin{split}{G_{f}}&=\|\widetilde{\nabla}f(x_{t})\|\frac{\left(h(x_{t}+r_{2}v_{f})-h(x_{t}-r_{2}v_{f})\right)}{2r_{2}},~~\textup{where }v_{f}=\frac{\widetilde{\nabla}f(x_{t})}{\|\widetilde{\nabla}f(x_{t})\|}\\ [{G_{h}}]_{:,i}&=\|\widetilde{\nabla}h_{i}(x_{t})\|\frac{\left(h(x_{t}+r_{2}v_{h,i})-h(x_{t}-r_{2}v_{h,i})\right)}{2r_{2}},~~\textup{where }v_{h,i}=\frac{\widetilde{\nabla}h_{i}(x_{t})}{\|\widetilde{\nabla}h_{i}(x_{t})\|},\end{split}$ (11)
Here $\widetilde{\nabla}h_{i}(x_{t})$ is the transpose of the $i$-th row of $\widetilde{J}_{h}(x_{t})$, i.e. $\widetilde{\nabla}h_{i}(x_{t})\!=\![\widetilde{J}_{h}(x_{t})]_{i,:}^{\top}$. Then, given $G_{f},G_{h}$, according to Eq. 10, we can set the Lagrangian multiplier $\lambda$ to be $\lambda_{t}=-(G_{h})^{-1}(G_{f}-Kh(x_{t}))$, which leads to our zeroth-order feedback linearization algorithm (ZOFL). The full procedure is summarized in Algorithm 1.

### v2

# Zeroth-Order Constrained Optimization from a Control Perspective R. Zhang, G. Zardini, A. Ozdaglar, J. Shamma, N. Li

@@ Abstract
Safe derivative-free optimization under unknown constraints is a fundamental challenge in modern learning and control. Existing zeroth-order (ZO) methods typically still assume access to a first-order oracle of the constraint functions or restrict attention to convex settings, leaving nonconvex optimization with black-box constraints largely unexplored. We propose the zeroth-order feedback-linearization (ZOFL) algorithm for ZO constrained optimization that enforces feasibility without access to the first-order oracle of the constraint functions and applies to both equality and inequality constraints. The proposed approach relies only on noisy, sample-based gradient estimates obtained via two-point estimators, yet provably guarantees constraint satisfaction under mild regularity conditions. It adopts a control-theoretic perspective on ZO constrained optimization and leverages feedback linearization, a nonlinear control technique, to enforce feasibility. Finite-time bounds on constraint violation and asymptotic global convergence guarantees are established for the ZOFL algorithm. A midpoint discretization variant is further developed to improve feasibility without sacrificing optimality. Empirical results demonstrate that ZOFL consistently outperforms standard ZO baselines, achieving competitive objective values while maintaining feasibility.
keywords:
zeroth-order optimization; constrained optimization; feedback linearization; nonlinear control; feasibility guarantees
{MSCcodes}
90C56, 93B52, 65K10
@@ 1 Introduction
Among derivative-free approaches, zeroth-order methods have attracted significant attention due to their simplicity and scalability to high dimensions [wang18stochastic, nesterov2017random]. The core idea is to build stochastic gradient estimators via finite differences of function evaluations and then plug them into standard gradient-based updates [berahas2022theoretical]. For instance, two-point estimators perturb the decision along random isotropic directions (Gaussian or uniform on the unit sphere) and combine the function values to approximate gradients [nesterov2017random, tang_zeroth-order_2023, ren2023escaping]. When combined with gradient descent, such estimators yield provable converge rates for unconstrained optimization.
In the constrained setting, most existing ZO algorithms assume black-box objectives but white-box constraints. Explicit knowledge of the constraint set enables efficient projections or local linearizations, ensuring feasibility. This has led to a variety of algorithms, including projection–gradient-descent [tang_zeroth-order_2023, liu_zeroth-order_2018, chen_zo-adamm_2019, ghadimi_mini-batch_2016], Frank–Wolfe–type methods [sahu_decentralized_2020, liu24zeroth-order, balasubramanian_zeroth-order_2019], and Sequential Quadratic Programming (SQP)-style approaches
@@
[curtis_almost-sure_2023, berahas2025sequential]. However, in many safe learning settings, such as safe RL or chance-constrained optimization [achiam2017constrained, cui2020adaptive, zhang2011chance], the constraint functions themselves are unknown. Here, only noisy zeroth-order estimates of the constraint gradients are available, and feasibility becomes much harder to enforce. Most existing work in this regime focuses on convex problems and relies on primal–dual schemes [zhou_zeroth-order_2025, chen_model-free_2022, yi_linear_2021, nguyen_stochastic_2022, li_zeroth-order_2022, maheshwari22, liu_min-max_2020, chen_continuous-time_2023, hu_gradient-free_2024]. For nonconvex problems, guarantees are scarce: some works provide only empirical evidence [zhou_zeroth-order_2025], while others require solving expensive convex subproblems at each step [nguyen_stochastic_2022, li_zeroth-order_2022]. Feasibility in these settings remains fragile, typically degrading as noise in the gradient estimators increases [oztoprak_constrained_2021].
@@
In contrast, in the first-order (FO) setting, SQP methods are known to be highly effective for nonconvex constrained optimization. They often outperform primal–dual approaches in practice, particularly when the number of constraints is small relative to the dimension of the decision variable [gould2003galahad, liu2018comparison, zhang2020proximal, applegate2024infeasibility]. Another complementary line of work interprets constrained optimization through a control-theoretic lens [cerone_new_2024, zhang2025constrainedoptimizationcontrolperspective]. Here, the optimization dynamics are viewed as a controlled system: the primal variables are states, the Lagrange multipliers are control inputs, and finding a first-order KKT point corresponds to steering the system to a feasible equilibrium. This perspective enables the use of nonlinear control tools, such as feedback linearization (FL), to design algorithms. Recent work shows that, under suitable conditions, FL-based schemes recover SQP-like updates and achieve strong performance on nonconvex problems [zhang2025constrainedoptimizationcontrolperspective].
@@
Despite their promise in first-order optimization, FL and SQP approaches have not been systematically studied in the zeroth-order regime. Extending them is nontrivial: FL and SQP depends critically on precise first-order information, but in the zeroth-order setting only noisy estimates are available, breaking the mechanisms that ensure feasibility. This raises the fundamental question:
@@
Our Contributions. We develop a zeroth-order constrained optimization framework that extends feedback linearization ideas to the derivative-free regime, inspired by control and dynamical systems perspective [cerone_new_2024, zhang2025constrainedoptimizationcontrolperspective]. First, we show how to construct an FL scheme tailored to dynamics evolving under noisy gradient information, in contrast to prior approaches that rely on convex relaxations or primal–dual surrogates. Second, we demonstrate that full Jacobians are unnecessary: it is enough to approximate a small set of Jacobian–vector products, which can be efficiently estimated via two-point zeroth-order queries. Third, we establish theoretical guarantees (Theorem 3.1 and Theorem 4.1) showing that constraint violations contract toward zero with high probability, up to controllable approximation and discretization errors. Finally, we provide empirical evidence that our method consistently achieves stronger feasibility performance than standard baselines, while achieving competitive objective values.
@@ 2 Preliminaries
We consider constrained optimization problems of the form
(1) $\min_{x\in\mathbb{R}^{n}}f(x)\quad\text{s.t. }h(x)=0,$
@@
(2) $\nabla f(x)+J_{h}(x)^{\top}\lambda=0,\quad h(x)=0.$
@@ 2.1 First-order Constrained Optimization: A Control Perspective
Recent works [cerone_new_2024, zhang2025constrainedoptimizationcontrolperspective] interpret constrained optimization from a control perspective, offering new insights and enabling novel algorithmic designs in the first-order optimization regime. The key idea is to reinterpret Eq. 2 as the equilibrium of a dynamical system. Specifically, define the updates
(3) $x_{t+1}-x_{t}=-\eta_{t}\!\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right)\!,~~~y_{t}=h(x_{t}),$
where $x_{t}$ is the system state, $y_{t}$ the constraint output, $\lambda_{t}$ the control input, and $\eta_{t}>0$ is the stepsize. Note that (3) incoorporates a wide range of optimization algorithms (cf. [nocedal_numerical_2006, neal2011distributed, chambolle2011first])
@@
(4) $x_{t+1}-x_{t}=-b(x_{t})+A(x_{t})\lambda_{t},$
by introducing a new input that cancels the nonlinearities [isidori1985nonlinear, henson1997feedback]. If $G(x)$ is invertible, one can apply a feedback transformation by introducing a new (virtual) control input $u_{t}$ and redefining the original input as  $\lambda_{t}=A(x_{t})^{-1}(b(x_{t})+u_{t})$. Substituting this expression into (4) yields: $x_{t+1}-x_{t}=u_{t}$, a linear system for which standard stabilizing controllers are available.
@@
(5) $y_{t+1}-y_{t}\approx J_{h}(x_{t})(x_{t+1}-x_{t})=-\underbrace{\eta_{t}J_{h}(x_{t})\nabla f(x_{t})}_{b(x_{t})}\underbrace{-\eta_{t}J_{h}(x_{t})J_{h}(x_{t})^{\top}}_{A(x_{t})}\lambda_{t},$
where the terms can be viewed as $b(x_{t})$ and $A(x_{t})\lambda_{t}$ in Eq. 4. Hence choosing
$\lambda_{t}=-\big(J_{h}(x_{t})J_{h}(x_{t})^{\top}\big)^{-1}\!\left(J_{h}(x_{t})\nabla f(x_{t})+u_{t}\right),$
cancels the nonlinear dependence and yields the linearized dynamics $y_{t+1}-y_{t}\approx\eta_{t}u_{t}$. In this paper, we consider a specific linear controller $u_{t}=-Ky_{t}$ where $K\in\mathbb{R}^{m\times m}$ is a positive definite matrix, thus we get $y_{t+1}-y_{t}\approx-Ky_{t}$ and hence the constraints converge exponentially to zero. This design gives the first-order feedback linearization (FO-FL) method:
FO-FL (Equality Constraints) [cerone_new_2024, zhang2025constrainedoptimizationcontrolperspective] (6) $\begin{split}x_{t+1}-x_{t}&=-\eta_{t}\left(\nabla f(x_{t})+J_{h}(x_{t})^{\top}\lambda_{t}\right),\\ \lambda_{t}&=-\left(J_{h}(x_{t})J_{h}(x_{t})^{\top}\right)^{-1}\!\left(J_{h}(x_{t})\nabla f(x_{t})-Kh(x_{t})\right).\end{split}$
FO-FL has been shown to effectively handle nonlinear dynamics, making it well-suited for nonconvex constrained optimization [cerone_new_2024, schropp_dynamical_2000, zhang2025constrainedoptimizationcontrolperspective].
@@ 2.2 Zeroth-order Constrained Optimization: Baseline and Challenges
The absence of first-order information motivates the use of stochastic finite-difference estimators for $\nabla f(x)$ and $J_{h}(x)$. A standard choice is the two-point estimator (cf. [nesterov2017random, ren2023escaping, tang_distributed_2021]):
(7) $\begin{split}\widetilde{\nabla}f(x_{t})&\!=\!\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}\frac{f(x_{t}\!+\!r_{1}u_{i})\!-\!f(x_{t}\!-\!r_{1}u_{i})}{2r_{1}}\,u_{i},\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\\ \widetilde{J}_{h}(x_{t})&\!=\!\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}\frac{h(x_{t}\!+\!r_{1}u_{i})\!-\!h(x_{t}\!-\!r_{1}u_{i})}{2r_{1}}\,u_{i}^{\top},\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\end{split}$
@@
ZO-baseline for Equality-Constrained Optimization (8) $\begin{split}x_{t+1}-x_{t}&=-\eta_{t}\!\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right),\\ \lambda_{t}&=-\!\left(\widetilde{J}_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\!\!\left(\widetilde{J}_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\right).\end{split}$
This approach has been explored in recent work on noisy or biased estimators [oztoprak_constrained_2021]. However, it suffers from a critical drawback: constraint satisfaction is no longer guaranteed. To see this, note that the constraint dynamics become
$\textstyle\qquad h(x_{t+1})-h(x_{t})\approx J_{h}(x_{t})(x_{t+1}-x_{t})$
@@ 3.1 Algorithm
(9) $\textstyle h(x_{t+1})\!-\!h(x_{t})\!\approx\!J_{h}(x_{t})(x_{t+1}\!-\!x_{t})\!=\!-\eta_{t}\!\left(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\!-\!J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right).$
@@
(10) $\textstyle\lambda_{t}=-\!\left(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\!\big(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})\big),$
@@
Insight. A closer examination reveals that full access to $J_{h}(x_{t})$ is unnecessary: it suffices to compute the Jacobian–vector products $J_{h}(x_{t})\,\widetilde{\nabla}f(x_{t})$ and $J_{h}(x_{t})\,\widetilde{J}_{h}(x_{t})^{\top}$. Equivalently, one only needs directional derivatives of $h$ along the directions $\widetilde{\nabla}f(x_{t})$ and the rows of $\widetilde{J}_{h}(x_{t})$, rather than the full Jacobian. Crucially, these Jacobian–vector products can be efficiently approximated using standard two-point estimators as follows, thereby rendering the scheme implementable in the ZO setting:
(11) $\begin{split}{G_{f}}&=\|\widetilde{\nabla}f(x_{t})\|\frac{\left(h(x_{t}+r_{2}v_{f})-h(x_{t}-r_{2}v_{f})\right)}{2r_{2}},~~\textup{where }v_{f}=\frac{\widetilde{\nabla}f(x_{t})}{\|\widetilde{\nabla}f(x_{t})\|}\\ [{G_{h}}]_{:,i}&\!=\!\|\widetilde{\nabla}h_{i}(x_{t})\|\frac{\left(h(x_{t}\!+\!r_{2}v_{h,i})\!-\!h(x_{t}\!-\!r_{2}v_{h,i})\right)}{2r_{2}},~\textup{where }v_{h,i}=\frac{\widetilde{\nabla}h_{i}(x_{t})}{\|\widetilde{\nabla}h_{i}(x_{t})\|},\end{split}$
Here $\widetilde{\nabla}h_{i}(x_{t})$ is the transpose of the $i$-th row of $\widetilde{J}_{h}(x_{t})$, i.e. $\widetilde{\nabla}h_{i}(x_{t})\!=\![\widetilde{J}_{h}(x_{t})]_{i,:}^{\top}$. The normalization factors $\|\widetilde{\nabla}f(x_{t})\|$ and $\|\widetilde{\nabla}h_{i}(x_{t})\|$ are introduced to convert directional finite differences along unit vectors into Jacobian–vector products along the original directions. Indeed, since $v_{f}$ and $v_{h,i}$ are unit-norm directions, the centered difference $\frac{h(x_{t}+r_{2}v)-h(x_{t}-r_{2}v)}{2r_{2}}$ approximates $J_{h}(x_{t})\,v$. Multiplying by the corresponding norm recovers $J_{h}(x_{t})\,\widetilde{\nabla}f(x_{t})$ and $J_{h}(x_{t})\,\widetilde{\nabla}h_{i}(x_{t})$, respectively. Thus, $G_{f}$} and $G_{h}$} are finite-difference estimates of the Jacobian–vector products $J_{h}(x_{t})\,\widetilde{\nabla}f(x_{t})$ and $J_{h}(x_{t})\,\widetilde{J}_{h}(x_{t})^{\top}$. Then, we can set the Lagrangian multiplier $\lambda$ to be $\lambda_{t}=-(G_{h})^{-1}(G_{f}-Kh(x_{t}))$, which leads to our zeroth-order feedback linearization algorithm (ZOFL). The full procedure is summarized in Algorithm 1.